OpenStudy (anonymous):

The fan guys games rental store charges an annual fee of $5 plus $5.50 per game rented. The game bank charges as annual fee of $17 plus $2.50 per game. For how many game rentals will the cost be the same at both stores? what is that cost?

OpenStudy (gw2011):

Let x = the number of games 5+5.50x=17+2.50x 5.50x-2.50x=17-5 3x=12 x=4 games 5+5.50(4)=17+2.50(4) 27=27 Therefore, $27 is the cost

OpenStudy (anonymous):

thank you very much (:

OpenStudy (gw2011):

You're welcome